OpenStudy (anonymous):

A group of friends are planning a road trip. If they all contribute $24, then they would be $6 short of the total cost of the trip. If they all contribute $25, then they would have $3 more than the total cost of the trip. What is the total cost of the trip?

OpenStudy (anonymous):

Let's call the number of friends \(\bf x\) and the cost of the road trip \(\bf y\). Then we can write out equations for the two situations given to us:\[\bf 24x=y-6\]\[\bf 25x=y+3\]You now have a system of equations through which you can solve for each variable. Can you do that? @depablo